# Confidence on the Focal: Conformal Prediction with Selection-Conditional Coverage

### v1

@@ B.5 Proof of Proposition 4
${\widetilde{T}}_{\textnormal{topk}}=\inf\Big{\{}t\in{\mathcal{T}}:\sum_{i\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{i}\leq t\}\geq m-K\Big{\}}.$
Abusing the notation a bit, we call $T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}$ and ${\widetilde{T}}_{\textnormal{topk}}^{\textnormal{swap}(i,j)}$ the corresponding stopping times defined with $({\mathcal{D}}_{\textnormal{calib}}^{\textnormal{swap}(i,j)}$, ${\mathcal{D}}_{\textnormal{test}}^{\textnormal{swap}(i,j)})$. Since ${\widetilde{T}}_{\textnormal{topk}}$ is agnostic to the swapping of unit $i$ and $n+j$, ${\widetilde{T}}^{\textnormal{swap}(i,j)}_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$.
We claim that when $S_{n+j}>T_{\textnormal{topk}}$, $T_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$, To see this, note first that for any $t\in{\mathcal{T}}$,
$\sum_{i\in{\mathcal{I}}_{\textnormal{test}}}\mathds{1}\{S_{i}\leq t\}\geq\sum_{i\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{i}\leq t\}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ {\widetilde{T}}_{\textnormal{topk}}\geq T_{\textnormal{topk}}.$
@@
$\sum_{i\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{i}\leq T_{\textnormal{topk}}\}\overset{\rm(a)}{=}\sum_{i\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{i}\leq T_{\textnormal{topk}}\}+\mathds{1}\{S_{n+j}\leq T_{\textnormal{topk}}\}\overset{\rm(b)}{\geq}m-K.$ (42)
@@
Similarly, we have ${\widetilde{T}}_{\textnormal{topk}}\geq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}$ by construction. When $S_{i}>T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}$,
$\sum_{\ell\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{\ell}\leq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\}=\sum_{\ell\in{\mathcal{I}}_{j}^{c}}\mathds{1}\{S_{\ell}\leq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\}+\mathds{1}\{S_{i}\leq{\widetilde{T}}_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\}\geq m-K.$
By the definition of $T_{\textnormal{topk}}$, we then have $T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\geq T_{\textnormal{topk}}$, and therefore $T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}=T_{\textnormal{topk}}$.
Fix $j\in\widehat{{\mathcal{S}}}_{\textnormal{topk}}$, we proceed to show that ^$\widehat{{\mathcal{R}}}_{\textnormal{fixed}}=\widehat{{\mathcal{R}}}_{n+j}$. For any $i\in\widehat{{\mathcal{R}}}_{\textnormal{topk}}$, there is $S_{i}>T_{\textnormal{topk}}$. We then have $S_{i}>T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}$, since $T_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}\geq T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}$. Therefore, $T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$, and
@@
wher step (a) follows from the fact that $S_{n+j}^{\textnormal{swap}(i,j)}=S_{i}$ and $S_{i}>{\widetilde{T}}_{\textnormal{topk}}$. As a result, $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$ and $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}\in\mathfrak{S}$. On the other hand, for any $i\in\widehat{{\mathcal{R}}}_{n+j}$, we have that $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$. Then there is
$S_{i}>T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}\Rightarrow S_{i}>{\widetilde{T}}_{\textnormal{topk}}\Rightarrow S_{i}>T_{\textnormal{topk}}.$ (43)
@@
B.6 Proof of Proposition 5
Fix $j\in\widehat{{\mathcal{S}}}$. Define the proxy stopping time that is agnostic to the swapping of units $i$ and $n+j$:
@@ B.6 Proof of Proposition 5
$\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}\leq{\widetilde{T}}_{\textnormal{cq}}\}=\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{i}\leq{\widetilde{T}}_{\textnormal{cq}}\}+\mathds{1}\{S_{n+j}\leq{\widetilde{T}}_{\textnormal{cq}}\}\geq qn,$ (44)
where the last inequality follows from the definition of ${\widetilde{T}}_{\textnormal{cq}}$. By the definition of $T_{\textnormal{cq}}$, (44) implies that ${\widetilde{T}}_{\textnormal{cq}}\geq T_{\textnormal{cq}}$ and therefore $T_{\textnormal{cq}}={\widetilde{T}}_{\textnormal{cq}}$.
Similarly, we have ${\widetilde{T}}_{\textnormal{cq}}\leq T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}$ by construction. When $S_{i}>{\widetilde{T}}_{\textnormal{cq}}$,
$\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{S_{\ell}\leq{\widetilde{T}}_{\textnormal{cq}}\}+\mathds{1}\{S_{n+j}\leq{\widetilde{T}}_{\textnormal{cq}}\}=\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{S_{\ell}\leq{\widetilde{T}}_{\textnormal{cq}}\}+\mathds{1}\{S_{n+j}\leq{\widetilde{T}}_{\textnormal{cq}}\}\geq qn,$
with the last step following from the definition of ${\widetilde{T}}_{\textnormal{cq}}$. We then have ${\widetilde{T}}_{\textnormal{cq}}\geq T_{\textnormal{cq}}^{\textnormal{swap}(i,j)}$ and therefore $T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}={\widetilde{T}}_{\textnormal{cq}}$.
@@
=$\overset{\rm(a)}{=}\{j\}\cup\{\ell\in[m]\backslash\{j\}:S_{n+\ell}>{\widetilde{T}}_{\textnormal{cq}}\}=\widehat{{\mathcal{S}}}_{\textnormal{cq}},$
@@
where step (a) follows from the fact that $S^{\textnormal{swap}(i,j)}_{n+j}=S_{i}$ and $S_{i}>{\widetilde{T}}_{\textnormal{cq}}$. The above implies that $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$ and and we have $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}\in\mathfrak{S}$. Reversely, for $i\in{\mathcal{I}}_{\textnormal{calib}}$ such that $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$,
$S_{n+j}^{\textnormal{swap}(i,j)}>T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}\Rightarrow S_{i}>{\widetilde{T}}_{\textnormal{cq}}\Rightarrow S_{i}>T_{\textnormal{cq}},$
completing the proof.
B.7 Proof of Proposition 6
To start, we claim that the $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}$ is equivalent to
${\widetilde{S}}_{\textnormal{fixed}}=\{j\in[m]:\widehat{{\mathcal{S}}}_{n+j}\geq T\},\text{ where }T=\inf\bigg{\{}t\in{\mathcal{T}}:1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq t,Y_{i}\leq c_{i}\}\geq q(n+1)\bigg{\}}.$
To see this, suppose $j\in\widehat{{\mathcal{S}}}_{\textnormal{fixed}}$, then $p_{j}\leq q$. We can check
$\frac{1}{n+1}\Big{(}1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+j},Y_{i}\leq c_{i}\}\Big{)}=p_{j}\leq q.$
By the definition of $T$, $\widehat{S}_{n+j}\geq T$. Conversely, if $\widehat{S}_{n+j}\geq T$, then
$p_{j}=\frac{1}{n+1}\Big{(}1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{n+j},Y_{i}\leq c_{i}\}\Big{)}\leq\Big{(}1+\sum_{i\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{i}\geq T,Y_{i}\leq c_{i}\}\Big{)}\leq q\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ j\in\widehat{{\mathcal{S}}}_{\textnormal{fixed}}.$
We then have $\widehat{{\mathcal{S}}}_{\textnormal{fixed}}={\widetilde{S}}_{\textnormal{fixed}}$. We shall work with ${\widetilde{S}}_{\textnormal{fixed}}$ in the following.
Fix $j\in{\widetilde{S}}_{\textnormal{fixed}}$ and a selection taxonamy $\mathfrak{S}$ such that ${\widetilde{S}}_{\textnormal{fixed}}\in\mathfrak{S}$. We prove the proposition for $y\leq c_{n+j}$ and $y>c_{n+j}$ separately. For notational simplicity, we define the swapped stopping time for any $i\in{\mathcal{I}}_{\textnormal{calib}}$ and $y\in{\mathcal{Y}}$ as
$T^{{\textnormal{swap}(i,j)}}(y)=\inf\bigg{\{}t\in{\mathcal{T}}:1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq t,Y_{i}\leq c_{i}\}+\mathds{1}\{\widehat{S}_{n+j}\geq t,y\leq c_{n+j}\}\leq q(n+1)\bigg{\}}.$
Then $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}(y)=\{\ell\in[m]:\widehat{S}^{\textnormal{swap}(i,j)}_{n+\ell}\geq T^{\textnormal{swap}(i,j)}(y)\}$
Case 1: $y\leq c_{n+j}$.
The goal is to show that ^$\widehat{{\mathcal{R}}}_{n+j}(y)=\widehat{{\mathcal{R}}}^{\textnormal{fixed}}_{n+j}(y)$. We prove this for $Y_{i}\leq c_{i}$ and $Y_{i}>c_{i}$.
  • $Y_{i}\leq c_{i}$.
    Suppose $i\in\widehat{{\mathcal{R}}}_{n+j}(y)$. Then $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$, i.e., $\widehat{S}_{i}\geq T^{\textnormal{swap}(i,j)}(y)$ and
    ${1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq\widehat{S}_{i},Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq\widehat{S}_{i},y\leq c_{n+j}\}}\leq q(n+1).$ (45)
    Since $y\leq c_{n+j}$, the above inequality is equivalent to
    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{S}_{\ell}\geq\widehat{S}_{i},Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq\widehat{S}_{i}\}\leq q(n+1).$ (46)
    Since $\mathds{1}\{\widehat{S}_{i}\geq\widehat{S}_{i},Y_{i}\leq c_{i}\}=1$, we further have
    ${\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}}\mathds{1}\{\widehat{S}_{\ell}\geq\widehat{S}_{i},Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq\widehat{S}_{i}\}}\leq q(n+1).$
    The above and the definition of $T_{j}^{\textnormal{fixed}}(1,0)$ implies that $\widehat{{\mathcal{S}}}_{i}\geq T_{j}^{\textnormal{fixed}}(1,0)$. Next, note that
    $1+\sum_{\ell\in{\mathcal{I}}_{\textnormal{calib}}\backslash\{i\}}\mathds{1}\{\widehat{{\mathcal{S}}}_{\ell}\geq T_{j}^{\textnormal{fixed}}(1,0),Y_{\ell}\leq c_{\ell}\}+\mathds{1}\{\widehat{S}_{n+j}\geq T_{j}^{\textnormal{fixed}}(1,0),y\leq c_{n+j}\}$

### v3

@@ D.7 Proof of Proposition 6
${\widetilde{T}}_{\textnormal{topk}}=\inf\Big{\{}t\in{\mathcal{T}}:\sum_{\ell\in[m]\backslash\{j\}}\mathds{1}\{S_{n+\ell}\leq t\}\geq m-K\Big{\}}.$
Abusing the notations a bit, we call $T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}$ and ${\widetilde{T}}_{\textnormal{topk}}^{\textnormal{swap}(i,j)}$ the corresponding stopping times defined with $({\mathcal{D}}_{\textnormal{calib}}^{\textnormal{swap}(i,j)}$, ${\mathcal{D}}_{\textnormal{test}}^{\textnormal{swap}(i,j)})$. By definition, it is straightfoward to check that ${\widetilde{T}}^{\textnormal{swap}(i,j)}_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$.
We claim that $S_{n+j}>T_{\textnormal{topk}}$ implies $T_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$. To see this, note first that for any $t\in{\mathcal{T}}$,
$\sum_{\ell\in[m]}\mathds{1}\{S_{n+\ell}\leq t\}\geq\sum_{\ell\in[m]\backslash\{j\}}\mathds{1}\{S_{n+\ell}\leq t\}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ {\widetilde{T}}_{\textnormal{topk}}\geq T_{\textnormal{topk}}.$
@@
$\sum_{\ell\in[m]\backslash\{j\}}\mathds{1}\{S_{n+\ell}\leq T_{\textnormal{topk}}\}\overset{\rm(a)}{=}\sum_{\ell\in[m]\backslash\{j\}}\mathds{1}\{S_{n+\ell}\leq T_{\textnormal{topk}}\}+\mathds{1}\{S_{n+j}\leq T_{\textnormal{topk}}\}\overset{\rm(b)}{\geq}m-K.$ (44)
@@
We can similarly show that $T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$ when $S_{i}>T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}$. By construction, we have ${\widetilde{T}}_{\textnormal{topk}}\geq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}$. Next,
$\sum_{\ell\in[m]\backslash\{j\}}\mathds{1}\{S_{n+\ell}\leq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\}=\sum_{\ell\in[m]\backslash\{j\}}\mathds{1}\{S_{n+\ell}\leq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\}+\mathds{1}\{S_{i}\leq T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\}\geq m-K.$
By the definition of ${\widetilde{T}}_{\textnormal{topk}}$, we then have $T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}\geq{\widetilde{T}}_{\textnormal{topk}}$, and therefore $T_{\textnormal{topk}}^{\textnormal{swap}(i,j)}={\widetilde{T}}_{\textnormal{topk}}$.
We now proceed to show that ^$\widehat{{\mathcal{R}}}_{\textnormal{fixed}}=\widehat{{\mathcal{R}}}_{n+j}$. For any $i\in\widehat{{\mathcal{R}}}_{\textnormal{topk}}$, it holds that $S_{i}>T_{\textnormal{topk}}$. Then since $T_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}\geq T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}$, we know that $S_{i}>T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}$. Therefore, $T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}$, and
@@
wher step (a) follows from the fact that $S_{n+j}^{\textnormal{swap}(i,j)}=S_{i}$ and $S_{i}>{\widetilde{T}}_{\textnormal{topk}}$. As a result, $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$ and $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}\in\mathfrak{S}$, i.e., $i\in\widehat{{\mathcal{R}}}_{n+j}$. On the other hand, for any $i\in\widehat{{\mathcal{R}}}_{n+j}$, it holds that $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$. Then,
$S_{i}>T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ T^{\textnormal{swap}(i,j)}_{\textnormal{topk}}={\widetilde{T}}_{\textnormal{topk}}=T_{\textnormal{topk}}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ S_{i}>T_{\textnormal{topk}}.$ (45)
@@
D.8 Proof of Proposition 7
Fix $j\in\widehat{{\mathcal{S}}}_{\textnormal{cq}}$. Define the proxy stopping time that is agnostic to the swapping of units $i$ and $n+j$:
@@ D.8 Proof of Proposition 7
$\sum_{i\in[n]}\mathds{1}\{S_{i}\leq{\widetilde{T}}_{\textnormal{cq}}\}=\sum_{i\in[n]}\mathds{1}\{S_{i}\leq{\widetilde{T}}_{\textnormal{cq}}\}+\mathds{1}\{S_{n+j}\leq{\widetilde{T}}_{\textnormal{cq}}\}\geq qn,$ (46)
where the first step is because $S_{n+j}>{\widetilde{T}}_{\textnormal{cq}}$ and the second step follows from the definition of ${\widetilde{T}}_{\textnormal{cq}}$. By the definition of $T_{\textnormal{cq}}$, (46) implies that ${\widetilde{T}}_{\textnormal{cq}}\geq T_{\textnormal{cq}}$ and therefore $T_{\textnormal{cq}}={\widetilde{T}}_{\textnormal{cq}}$.
Similarly, we claim that $T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}={\widetilde{T}}_{\textnormal{cq}}$ when $S_{i}>{\widetilde{T}}_{\textnormal{cq}}$. By construction, we have ${\widetilde{T}}_{\textnormal{cq}}\leq T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}$. When $S_{i}>{\widetilde{T}}_{\textnormal{cq}}$,
$\sum_{\ell\in[n]\backslash\{i\}}\mathds{1}\{S_{\ell}\leq{\widetilde{T}}_{\textnormal{cq}}\}+\mathds{1}\{S_{n+j}\leq{\widetilde{T}}_{\textnormal{cq}}\}=\sum_{\ell\in[n]}\mathds{1}\{S_{\ell}\leq{\widetilde{T}}_{\textnormal{cq}}\}+\mathds{1}\{S_{n+j}\leq{\widetilde{T}}_{\textnormal{cq}}\}\geq qn,$
with the first step following from $S_{i}>{\widetilde{T}}_{\textnormal{cq}}$ and the second step from the definition of ${\widetilde{T}}_{\textnormal{cq}}$. We then have ${\widetilde{T}}_{\textnormal{cq}}\geq T_{\textnormal{cq}}^{\textnormal{swap}(i,j)}$ and therefore $T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}={\widetilde{T}}_{\textnormal{cq}}$.
@@
=$\overset{\rm(a)}{=}\{j\}\cup\{\ell\in[m]\backslash\{j\}:S_{n+\ell}>T_{\textnormal{cq}}\}=\widehat{{\mathcal{S}}}_{\textnormal{cq}},$
@@
where step (a) follows from the fact that $S^{\textnormal{swap}(i,j)}_{n+j}=S_{i}$ and $S_{i}>{\widetilde{T}}_{\textnormal{cq}}$. The above implies that $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$ and that $\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}\in\mathfrak{S}$. Conversely, for $i\in[n]$ such that $j\in\widehat{{\mathcal{S}}}^{\textnormal{swap}(i,j)}$,
$S_{i}>T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ T^{\textnormal{swap}(i,j)}_{\textnormal{cq}}={\widetilde{T}}_{\textnormal{cq}}=T_{\textnormal{cq}}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ S_{i}>T_{\textnormal{cq}},$
which completes the proof.
Appendix E Auxiliary results
E.1 Connection between selection-conditional coverage and FCR control
Lemma 1 establishes the asympototic equivalence between the selection-conditional coverage and the false coverage rate (FCR) when the selection rule is asymptotically close to a threshold function.
Lemma 1.
Suppose $\{Z_{i}\}_{i=1}^{m+n}$ are i.i.d., and there exists some fixed function $f:{\mathcal{X}}\mapsto\mathbb{R}$ and fixed constant $\tau\in\mathbb{R}$ such that the selection set $\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ obeys $|\widehat{\mathcal{S}}\Delta\widehat{\mathcal{S}}^{*}|/|\widehat{\mathcal{S}}|{\to}0$ in probability as $n,m\rightarrow\infty$, where $\widehat{\mathcal{S}}^{*}=\{j\colon f(X_{n+j})\geq\tau\}$ and $\mathbb{P}(f(X_{n+1})\geq\tau)>0$. Assume that $\widehat{S}$ is equivariant to the permutation of the units in ${\mathcal{D}}_{\textnormal{test}}$ and that the prediction set takes the form $\widehat{C}_{\alpha,n+j}=\{y\colon V(X_{n+j},y)\leq\widehat{q}_{m,n}\}$, where $\widehat{q}_{m,n}\stackrel{{\scriptstyle\textnormal{p}}}{{\to}}q^{*}\in\mathbb{R}$, as $m,n\to\infty$. Assume also that the function $G(t):=\mathbb{P}(f(X_{n+1})\geq\tau,V_{n+1}>t)$ is continuous in $t$. Then,